# Faster, Cheaper, Better: Multi-Objective Hyperparameter Optimization for LLM and RAG Systems

### v1

@@ 1 Introduction
Multi-objective HO over a RAG pipeline is particularly challenging for several reasons. First, RAG pipelines naturally have a high number of parameters, leading to a large solution space. We identify at least 15 example choices, parameters, and hyperparameters in Table 1. Even if one has just a handful of possible values for each choice, the parameter space becomes intractably large for simple algorithms like grid search.  [^1] https://scikit-learn.org/stable/modules/generated/sklearn.model_selection.GridSearchCV.html Second, evaluating a RAG pipeline during the HO process is costly, with respect to both compute resources and time: it requires running the RAG system over multiple queries (where each iteration is bounded by the per-token inference time of the LLM), and then evaluating each output. Third, the objective evaluations can be noisy. The true characteristic of a RAG system cannot be computed directly, and requires sampling the evaluations from many queries. Relatedly, since in most cases LLMs are non-deterministic, combinations of hyperparameters need to be tested over multiple seeds.
@@
    We introduce an approach for multi-objective optimization over a unique set of hyperparameters of a RAG pipeline, including choices for the LLM and embedding models themselves. Our approach implements a single composite objective value called the hypervolume improvement (Guerreiro et al., 2021), and uses Bayesian optimization with an acquisition function that allows for an arbitrary number of noisy objective functions (qLogNEHVI)  (Daulton et al., 2021; Ament et al., 2023) to find the best RAG pipeline configuration.
@@ 3 Preliminaries and problem formulation
$\max_{\mathcal{\hat{P}}}\mathcal{HV}(\mathcal{\hat{P}}|\mathbf{r})$ (3)
@@ 4.2 Objectives
In this work, we use the term “safety” to refer to hallucination risks, or the risk that a RAG pipeline will return false information to the user. Hallucinations can cause significant downstream harm, particularly in high-stakes domains such as healthcare. In our experiments, we evaluate safety using the faithfulness metric defined in the Trustwise API.555 Trustwise documentation available at: https://trustwise.ai/docs Like previous work (Min et al., 2023; Es et al., 2023), faithfulness detects hallucinations by evaluating whether or not the response from a RAG system is supported by the context. The response is split into individual, “atomic” claims that are verified with respect to the context. Scores of these verifications are then aggregated into a single faithfulness score between 0 and 100 for each response, where 100 represents a completely “safe” response with no hallucinations. [^6] For those aiming to replicate our approach, there are open-source alternatives for evaluating safety that could be used instead, such as LlamaGuard (Inan et al., 2023).
@@ 5.2 Baselines
For our tasks, we lack access to a “ground truth” set of Pareto optimal configurations because we cannot evaluate the objective functions directly, and grid search is computationally unfeasible. Hence we use three baseline approaches that are applicable in MOO settings. The first is uniform sampling, which generates configurations independently, where each configuration from the solution space is equally probable. The second is Sobol sampling. Like latent hypercube sampling (Loh, 1996), this method generates configurations that guarantee good high-dimensional uniformity. It is commonly used to initialize optimization algorithms, including BO, and represents a sensible alternative to grid-search. Finally, we use qLogEHVI BO, the noiseless variant of our chosen acquisition function (Daulton et al., 2020).

### v2

@@ 1 Introduction
Multi-objective HO over a RAG pipeline is particularly challenging for several reasons. First, RAG pipelines naturally have a high number of parameters, leading to a large solution space. We identify at least 15 example choices, parameters, and hyperparameters in Table 1. Even if one has just a handful of possible values for each choice, the parameter space becomes intractably large for simple algorithms like grid search (Bergstra & Bengio, 2012). Second, evaluating a RAG pipeline during the HO process is costly, with respect to both compute resources and time: it requires running the RAG system over multiple queries (where each iteration is bounded by the per-token inference time of the LLM), and then evaluating each output. Third, the objective evaluations can be noisy. The true characteristic of a RAG system cannot be computed directly, and requires sampling the evaluations from many queries. Relatedly, since in most cases LLMs are non-deterministic, combinations of hyperparameters need to be tested over multiple seeds.
@@
    We introduce an approach for multi-objective optimization over a unique set of hyperparameters of a RAG pipeline, including choices for the LLM and embedding models themselves. Our approach implements a single composite objective value called the hypervolume indicator (Guerreiro et al., 2021), and uses Bayesian optimization with an acquisition function that allows for an arbitrary number of noisy objective functions (qLogNEHVI)  (Daulton et al., 2021; Ament et al., 2023) to find the best RAG pipeline configuration.
@@ 3 Preliminaries and problem formulation
$\operatorname*{arg\,max}_{\mathcal{\hat{P}}}\mathcal{HV}(\mathcal{\hat{P}}|\mathbf{r})$ (3)
@@ 4.2 Objectives
In this work, we use the term “safety” to refer to hallucination risks, or the risk that a RAG pipeline will return false information to the user. Hallucinations can cause significant downstream harm, particularly in high-stakes domains such as healthcare. In our experiments, we evaluate safety using the faithfulness metric defined in the Trustwise API.444 Trustwise documentation available at: https://trustwise.ai/docs Like previous work (Min et al., 2023; Es et al., 2023), faithfulness detects hallucinations by evaluating whether or not the response from a RAG system is supported by the context. The response is split into individual, “atomic” claims that are verified with respect to the context. Scores of these verifications are then aggregated into a single faithfulness score between 0 and 100 for each response, where 100 represents a completely “safe” response with no hallucinations. [^5] For those aiming to replicate our approach, there are open-source alternatives for evaluating safety that could be used instead, such as LlamaGuard (Inan et al., 2023).
@@ 5.2 Baselines
For our tasks, we lack access to a “ground truth” set of Pareto optimal configurations because we cannot evaluate the objective functions directly, and grid search is computationally unfeasible. Hence we use three baseline approaches that are applicable in MOO settings. The first is uniform sampling, which generates configurations independently, where each configuration from the solution space is equally probable. The second is Sobol sampling. Like Latin hypercube sampling (Loh, 1996), this method generates configurations that guarantee good high-dimensional uniformity. It is commonly used to initialize optimization algorithms, including BO, and represents a sensible alternative to grid-search. Finally, we use qLogEHVI BO, the noiseless variant of our chosen acquisition function (Daulton et al., 2020).